Six pears and three apples cost $3.90. Two pears and five apples cost $3.30. How much does one pear cost?
Hint: You have two unknowns (variables), so you need two equations. Write one equation for each situation and solve the system of equations.

Answer by stanbon(48559)

Six pears and three apples cost $3.90. Two pears and five apples cost $3.30. How much does one pear cost?
-----------
Let "p" be number of pears and "a" be number of apples
EQUATIONS:
1st: 6p+3a=3.90
2nd: 2p+5a=3.30
Multiply 2nd by 3 to get:
3rd: 6p+15a=9.90
Subtract 1st from 3rd to get:
4th: 12a=6
a=1/2 (apples cost 50 cents)
Substitute this value of a into 1st or into 2nd to solve for p:
2p+5(1/2)=3.3
2p+2.5=3.3
2p=0.8
p=0.4 (peaches cost 40 cents)
--------
Cheers,
Stan H.